### v2

@@ Abstract
We present a novel approach to the automated semantic analysis of legal texts using large language models (LLMs), targeting their transformation into formal representations in Defeasible Deontic Logic (DDL). We propose a structured pipeline that segments complex normative language into atomic snippets, extracts deontic rules, and evaluates them for syntactic and semantic coherence. Our methodology is evaluated across various LLM configurations, including prompt engineering strategies, fine-tuned models, and multi-stage pipelines, focusing on legal norms from the Australian Telecommunications Consumer Protections Code. Empirical results demonstrate promising alignment between machine-generated and expert-crafted formalizations, showing that LLMs — particularly when prompted effectively — can significantly contribute to scalable legal informatics.
@@ 1 Introduction
The idea of automated legal reasoning has been one of the cornerstones of AI and Law for a long time, with many concrete attempts since the seminal paper by Sergot and Kowalski [1] on the formal representation of the British Nationality Act (see [2, 3] for an overview of some of the most influential approaches). A recent OECD report [4] outlines the benefits of the adoption of automated legal reasoning and encoding legal provisions in a format that is processable by machines. The major obstacle to this vision is the knowledge representation bottleneck. Anecdotal data from many large-scale encoding projects suggests that an experienced coder can only encode 4 to 5 pages per day, with serious burnout concerns (a recent empirical experiment on legal coding confirms the rate of encoding [5]). The issue is further exacerbated by the proliferation of legal information and the rising complexity of regulatory environments, which have intensified the need for automated tools that can interpret and formalize normative documents. Accordingly, there is a need to have tools that can help with the encoding of legal instruments.
The idea of using NLP techniques, specifically categorial grammar-based approaches, to encode norms was advanced by [6]. Then it has been extended in [7] where it was tested on small-scale examples. Successively, it was adopted in [8] that introduced a manually supervised pipeline for extracting a formal representation, relying on deterministic parsing rules. While this approach resulted in reasonable outcomes, a successful extraction required many iterations and was sensitive to the specific format of the input. Furthermore, it did not lead to a significant reduction in the time needed to create a complete and fully functional encoding. On the other hand, [9] explored the use of ML-based NLP techniques for the normative encoding process. The key findings were that these approaches required very extensive training data (which was and still is not available), and that the performance was not comparable to the rule-based approach. More recent efforts have incorporated neural methods for legal information retrieval and summarization, but few have addressed the formalization task with the granularity and formal logic representation at the level we propose in this paper. For instance, the recent work [10] focuses on a single article from the Council Framework Decision 2002/584/JHA (European Arrest Warrant).
@@
In recent years, Large Language Models (LLMs) have emerged as powerful tools for understanding and generating natural language. However, their application to legal texts remains underexplored, particularly in tasks requiring semantic precision, such as the conversion of legal norms into machine-interpretable representations.
@@
This paper explores the feasibility and effectiveness of using LLMs to translate legal language into a formal representation. More specifically, we compare with the approach of [8], and we encode legal provisions in Defeasible Deontic Logic (DDL), a computational logic framework designed to reason about obligations, permissions, and prohibitions. The target corpus is the Australian Telecommunications Consumer Protections Code (TCP Code), characterized by complex, hierarchical rules. This dataset was also used in [8], and we compare our results with those reported there.
@@
Our central hypothesis is that, with suitable prompting and architectural configurations, LLMs can assist in extracting semantically valid and logically coherent deontic rules from unstructured legal text. The novel contribution of this work lies in the integration of prompt engineering techniques, evaluation metrics grounded in logical correctness, and comparative studies of different LLM architectures and training strategies.
@@
The remainder of this paper is structured as follows. Section 2 provides the necessary background on LLMs and the formal representation language DDL used in this study. Section 3 outlines the methodology, including the segmentation of legal texts into individual law snippets, their transformation into DDL, and the evaluation approach. Section 4 presents the experimental results, covering prompt engineering, multi-snippet processing strategies, fine-tuning, and two-stage pipelines. This section also includes a detailed comparison with the evaluation framework proposed by Dragoni et al. [8]. Section 5 discusses key limitations, such as challenges in legal implementation, handling of inter-snippet references, and atom reuse. Finally, Section 6 concludes the paper and outlines directions for future research.
@@ 2 Background
Defeasible Deontic Logic [11, 12] is a flexible and efficient rule-based non-monotonic formalism for the representation of legal norms and legal reasoning. The logic combines features of Defeasible Logic for the natural modeling of exceptions and defeasibility with concepts from Deontic Logic (i.e., obligations, permission, prohibition, compensatory obligations). A rule in DDL has the form
@@
LLMs can be broadly divided into two categories: traditional LLMs and reasoning LLMs. Traditional LLMs, such as the GPT series by OpenAI, DeepSeek-V3 or similar models, are primarily optimized for fluent language generation and general-purpose tasks. Their strength lies in producing coherent and contextually appropriate text. However, their capabilities in logical reasoning and structured problem-solving are limited. Reasoning LLMs represent a newer generation of models that are explicitly designed to perform structured reasoning tasks more effectively. These models consider additional training objectives, architectural innovations, or fine-tuning procedures that enhance their ability to perform logical inference, complex decision-making, and consistent multi-step problem-solving. Reasoning LLMs aim not only at linguistic fluency but also at improved logical accuracy and reliability in formal contexts.
@@
In this study, we consider models from both categories: (i) Traditional LLMs: GPT-4o, GPT-4o mini, DeepSeek-V3, and (ii) Reasoning LLMS: OpenAI o3, OpenAI o1, OpenAI o4-mini, OpenAI o3-mini, DeepSeek-R1. These models will be evaluated and compared based on their performance in formalizing legal texts into DDL representations.
@@ 3 Methodology
    Controls randomness; higher values yield more diverse outputs [13]. Set to 0 to prioritize consistency over creativity [14].
@@
    Governs nucleus sampling [15]; left at the default to avoid compounding effects with temperature, per API guidelines [16, 17].
@@
    Penalizes repeated tokens by frequency [18]. Disabled to ensure consistent terminology use.
@@
    Penalizes repeated tokens after first occurrence [18]. Also disabled for terminological consistency.
These settings were applied both during legal text segmentation and formalization into DDL. For OpenAI’s reasoning models, which do not accept these parameters, the reasoning_effort option was set to high.
@@
  • Fine-tuning of GPT-4o using a limited dataset of annotated examples to enhance task-specific performance.
  • Two-stage pipelines, using separate LLMs for atom extraction and rule generation to enhance consistency and limit error propagation
3.1 Segmentation into Law Snippets
Legal texts are initially segmented into manageable ”law snippets“ using DeepSeek R1. Enumerations in legal provisions are split into individual rules where appropriate, aiming at a balance between contextual completeness and token constraints.
A key challenge in instructing the LLMs was determining the optimal length for law snippets: overly long snippets risked losing critical information during formalization, whereas overly short ones hindered atom reuse. To address this, we instructed the model to split enumerations containing more than two elements into separate law snippets while preserving shorter ones intact.
3.2 Transformation into DDL
Each law snippet is transformed into DDL rules via various prompting strategies. These include Chain-Of-Instructions (CoI) prompting and few-shot learning using prompt variants with progressively enhanced instructions. We also evaluate a pipeline approach where atom extraction and DDL rule generation are handled by different LLMs in sequence.
Note that despite following OpenAI’s guidelines for achieving reproducible outputs [19], e.g., fixing the seed and temperature parameters, we observed non-deterministic behavior. This phenomenon can be attributed to inherent LLM stochasticity [20].
@@ 3.3 Evaluation
Consider, for instance, the following formalization of law snippet 8.2.1.a.xiv:
@@
This initial check is crucial to prevent the LLM from achieving a high score merely by formalizing only the simplest aspects of a problem.
@@
closeComplaint(X), -consent(X), -clausesCDEComplied(X) $\Rightarrow$
    [O] closeComplaint(X)
Note that the consequence of this rule also appears as its antecedent. However, this issue was later resolved by adding a corresponding instruction to the prompt.
@@
However, there are also more subtle issues filtered by this question, for example, when a LLM combines several aspects with a logical “and”, even though they should be connected with a logical “or” according to the legal text. This question also verifies whether formalizations that are not syntactically identical to another rule convey the same meaning and are therefore redundant.
@@
-consentConsumer(X), -compliedWithClauseC(X),
-compliedWithClauseD(X), -compliedWithClauseE(X) $\Rightarrow$
    [F] closeComplaint(X)
@@
In the success score calculation, we represent the outcomes of questions Q1 through Q6 using binary values: 1 for true (satisfied) and 0 for false (not satisfied).
We define $\text{Qi}(r)\in\{0,1\}$ as the evaluation of question Qi on rule $r$. For a given law snippet $l$, let $\mathcal{R}(l)$ denote the set of rules generated from $l$. We first introduce a modifier function $m(l)$, which reduces the success score by half if Q1 is not satisfied over the entire snippet:
$m(l)=\begin{cases}0.5&\text{if }\text{Q1}(l)=0,\\ 1&\text{if }\text{Q1}(l)=1,\end{cases}$
where $\text{Q1}(l)$ evaluates the overall satisfaction of Q1 across the law snippet $l$ (i.e., by considering all generated rules together). The success score $S(l)$ for an individual law snippet $l$ is then defined as:
$S(l)=m(l)\times\frac{1}{|\mathcal{R}(l)|}\sum_{r\in\mathcal{R}(l)}\frac{1}{5}\sum_{i=2}^{6}\text{Qi}(r).$
@@ 4 Results
First, we start with an initial experiment, where the LLMs are given a prompt with detailed instructions how to solve the problem, and the respective law snippet to formalize. We also perform different variations of the experiment, including different output formats and varied prompts. We then evaluate fine-tuned models. Finally, we implement a pipeline where two LLMs work together to solve the task. Specifically, one LLM is responsible for extracting the atoms and another for the actual formalization of the DDL rules.
In each of the following experiments, we provided LLMs with a prompt containing step-by-step instructions to guide the model through the extraction process. This approach is called Chain-of-Instructions (CoI) prompting [21]. Hence, the model is encouraged to solve each subtask step by step until the final answer is reached [22]. This method contrasts with Chain-of-Thoughts (CoT), which usually depends more on implicit reasoning [21] – especially for Zero-Shot-CoT, where just a sentence like “Let’s think step by step” is appended to the prompt [23].
Moreover, we use few-shot learning [24], where we provide the LLM with a few examples (input-output pairs) in the prompt to demonstrate how to solve the task.

### v3

@@ Abstract
We present a comprehensive approach to the automated formalization of legal texts using large language models (LLMs), targeting their transformation into Defeasible Deontic Logic (DDL). Our method employs a structured pipeline that segments complex normative language into atomic snippets, extracts deontic rules, and evaluates them for syntactic and semantic coherence. We introduce a refined success metric that more precisely captures the completeness of formalizations, and a novel two-stage pipeline with a dedicated refinement step to improve logical consistency and coverage. The evaluation procedure has been strengthened with stricter error assessment, and we provide comparative results across multiple LLM configurations, including newly released models and various prompting and fine-tuning strategies. Experiments on legal norms from the Australian Telecommunications Consumer Protections Code demonstrate that, when guided effectively, LLMs can produce formalizations that align closely with expert-crafted representations, underscoring their potential for scalable legal informatics.
@@ 1 Introduction
Automated legal reasoning has been a major endeavour of research in AI and Law. Numerous efforts have followed the seminal paper by [BNA] on formally representing the British Nationality Act in Prolog. [DBLP:journals/ail/Bench-CaponAAABBBBCFGGLLLMPSSTTVWW12] and [ail:30years] provide an overview of key approaches in this field. More recently, the idea of Rules as Code emerged as a general approach to speed up the development of law-based information systems to provide services to the citizens and businesses. The white paper by OECD by [MohunRobertsOECD2020] emphasises the advantages of adopting automated legal reasoning and encoding legal provisions in a machine-readable format. However, the main challenge to this vision lies in the knowledge representation bottleneck. Anecdotal evidence from numerous large-scale encoding projects suggests that an experienced coder can only encode about 4 to 5 pages per day. A recent empirical study by [Cristani:encoding] on legal coding confirms this rate. Furthermore, manually encoding large regulatory frameworks raises serious concerns about burnout. While a parallel encoding by a team of human coders seems a solution, an experiment by [ail2022:Alice] points out that, very likely, the parts encoded by different members of the team are incompatible with each other and reconciling them requires some considerable overhead.
The knowledge acquisition problem is exacerbated by the growing volume of legal information and the increasing complexity of regulatory environments. This has heightened the need for automated tools capable of interpreting and formalizing normative documents. Consequently, there is a pressing need for tools that can assist with encoding legal instruments.
The idea of employing NLP techniques, particularly categorical grammar-based approaches, for encoding norms was advanced by [WynerPeters]. This idea was further explored in [DBLP:conf/ruleml/WynerG13], where it was tested on small-scale examples. Subsequently, [DBLP:conf/aicol/DragoniVRG17] introduced a manually supervised pipeline for extracting a formal representation. This approach relied on deterministic parsing rules and yielded reasonable outcomes. However, successful extraction required numerous iterations and was susceptible to the specific format of the input. Furthermore, it did not result in a substantial reduction in the time required to create a comprehensive and fully functional encoding.
On the other hand, [jurisin19:nlp] explored the use of ML-based NLP techniques for the normative encoding process. Their key findings were that these approaches required extensive training data, which was and still is not readily available. Furthermore, their performance was not comparable to the approach proposed by [DBLP:conf/aicol/DragoniVRG17].
While recent efforts have incorporated neural methods for legal information retrieval and summarisation, few have addressed the formalisation task with the granularity and formal logic representation we propose in this paper. For example, the recent work [BilliPS24] focusses on a single article from the Council Framework Decision 2002/584/JHA (European Arrest Warrant). The approach by [icail25:sartor] proposes a pipeline based on LLM and Logical English (a structured controlled language) to translate the UK Highway Code into Prolog. In their experiment, they considered four rules from the code, and the LLM correctly translated three of them.
@@
Large Language Models (LLMs) have emerged as powerful tools for understanding and generating natural language. However, their application to legal texts remains underexplored, particularly in tasks requiring semantic precision, such as converting legal norms into machine-interpretable representations.
@@
This paper explores the feasibility and effectiveness of using LLMs to translate legal language into a formal representation. We compare this approach with the one proposed in [DBLP:conf/aicol/DragoniVRG17], where legal provisions are encoded in Defeasible Deontic Logic (DDL), a computational logic framework designed to reason about obligations, permissions, and prohibitions.
The target corpus is the Australian Telecommunications Consumer Protections Code (TCP Code), containing complex and hierarchical rules. This dataset was also used in [DBLP:conf/aicol/DragoniVRG17], and we compare our results with those reported there.
@@
Our central hypothesis is that, with suitable prompting and architectural configurations, LLMs can assist in extracting semantically valid and logically coherent deontic rules from unstructured legal text. The novel contribution of this work lies in the integration of prompt engineering techniques, evaluation metrics grounded in logical correctness, and comparative studies of different LLM architectures and training strategies. This work substantially extends [Horner2025]. Notable improvements include a revised formula for the success score that more precisely reflects the extent to which all aspects of a law text have been formalized by a LLM, evaluations of newly released models and the introduction of a two-stage pipeline with a refinement step. Additionally, the evaluation procedure has been refined through stricter error assessment.
@@
The remainder of this paper is structured as follows. Section 2 provides the necessary background on LLMs and the formal representation language DDL used in this study. Section 3 outlines the methodology, including the segmentation of legal texts into individual law snippets, their transformation into DDL, and the evaluation approach. Section 4 presents the experimental results, covering prompt engineering, multi-snippet processing strategies, fine-tuning, and two-stage pipelines. This section also includes a detailed comparison with the evaluation framework proposed by [DBLP:conf/aicol/DragoniVRG17]. Section 5 discusses key limitations, such as challenges in legal implementation and handling of inter-snippet references. Finally, Section 6 concludes the paper and outlines directions for future research.
@@ 2 Background
Defeasible Deontic Logic [jpl:permission, handbook:deontic] is a flexible and efficient rule-based non-monotonic formalism for the representation of legal norms and legal reasoning. The logic combines features of Defeasible Logic for the natural modeling of exceptions and defeasibility with concepts from Deontic Logic (i.e., obligations, permission, prohibition, compensatory obligations). A rule in DDL has the form
@@
LLMs can be broadly divided into two categories: traditional LLMs and reasoning LLMs. Traditional LLMs, such as the GPT series by OpenAI, DeepSeek-V3 or similar models, are primarily optimized for fluent language generation and general-purpose tasks. Their strength lies in producing coherent and contextually appropriate text. However, their capabilities in logical reasoning and structured problem-solving are limited. Reasoning LLMs represent a newer generation of models that are explicitly designed to perform structured reasoning tasks more effectively. These models consider additional training objectives, architectural innovations, or fine-tuning procedures that enhance their ability to perform logical inference, complex decision-making, and consistent multi-step problem-solving. Reasoning LLMs aim not only at linguistic fluency, but also at improved logical accuracy and reliability in formal contexts.
@@
In this study, we consider models from both categories: (i) Traditional LLMs: GPT-4.1, GPT-4o, GPT-4o mini, DeepSeek-V3 (both 0324 and the original version), Claude Sonnet 4 (without extended thinking), Llama 4 Maverick; and (ii) Reasoning LLMs: OpenAI o3, OpenAI o1, OpenAI o4-mini, OpenAI o3-mini, DeepSeek-R1 (both 0528 and the original version), Claude Sonnet 4 (with extended thinking). Out of these considered models, all models of DeepSeek as well as Llama 4 Maverick are open-source. These models will be evaluated and compared based on their performance in formalizing legal texts into DDL representations.
@@ 3 Methodology
    The temperature parameter controls the degree of randomness in the generation process. Higher values lead to more diverse outputs [DBLP:conf/icccrea/PeeperkornK0J24], facilitating exploratory or creative behavior [DBLP:conf/ccnlg/ManjavacasKBK17]. Therefore, temperature was set to 0.
@@
    Top-p sampling, also known as nucleus sampling, limits the pool of candidate tokens to those comprising the top cumulative probability $p$ [DBLP:conf/iclr/HoltzmanBDFC20]. A value of $p=1$ effectively disables nucleus sampling, allowing all tokens to be considered. The parameter was set at its default value of $1$, following recommendations from both OpenAI and DeepSeek that advise adjusting either temperature or top-p, but not both simultaneously [OpenAiApiReference, DeepSeekApiReference].
@@
    This parameter discourages the repeated generation of tokens that have already appeared by reducing their probability proportionally to their frequency [DBLP:journals/corr/abs-2408-10577]. While beneficial for enhancing lexical diversity, it is unsuitable for the present task, where consistent use of terminology across similar legal contexts is required. Therefore, the parameter was left at its default setting of 0.
@@
    The presence penalty penalizes the generation of tokens that have already been produced at least once, independently of their frequency [DBLP:journals/corr/abs-2408-10577]. Its purpose is to encourage novel token introduction. However, similar to the frequency penalty, consistent lexical choice is essential for reliable formalization, so this parameter was kept at its default value of 0.
These settings were applied both during legal text segmentation and formalization into DDL. For OpenAI’s reasoning models, which do not accept these parameters, the reasoning_effort option was set to high. Also for Claude Sonnet 4 with extended thinking enabled, these options were not available.
3.1 Segmentation into Law Snippets
Legal texts are initially segmented into manageable ”law snippets“ using DeepSeek-R1. A key challenge in instructing the LLMs was determining the optimal length for law snippets: overly long snippets risked losing critical information during formalization, whereas overly short ones hindered atom reuse. To address this, we instructed the model to split enumerations containing more than two elements into separate law snippets while preserving shorter ones intact.
3.2 Transformation into DDL
@@
  • Fine-tuning of GPT-4.1 and GPT-4o using a limited dataset of annotated examples to enhance task-specific performance.
  • Two-stage pipelines, where we explored two approaches: (i) using a separate LLM for atom extraction, and (ii) introducing a refinement step after the initial generation.
Despite adhering to the guidelines for achieving reproducible outputs by [OpenAiReproducibleOutputsReference], e.g., fixing the seed and temperature parameters (when supported by the model), we encountered non-deterministic behavior. This phenomenon can be attributed to inherent LLM stochasticity [blackwell2024]. Note that this behavior was also observed in the open-source models considered in this study.
@@ 3.3 Evaluation
Consider, for instance, the following formalization of law snippet 8.2.1.a.xiv, which states that a supplier may close a complaint only with the consumer’s consent or in compliance with clauses 8.2.1(c), (d), or (e).
@@
This initial check is crucial to prevent the LLM from achieving a high score merely by formalizing only the simplest aspects of a problem. Unlike subsequent dimension evaluations, however, Q1 is evaluated once per law snippet, not once per generated rule. Another important distinction between Q1 and the other checks is that the answer in this case is not a binary true/false value. Instead, it is expressed as a percentage indicating the proportion of gold standard rules for the respective law snippet that were attempted to be formalized. A rule is considered attempted if it can be meaningfully matched to a gold standard rule, even if the generated version contains minor flaws (e.g., suboptimal atom naming, redundant preconditions). The key point is that Q1 measures attempted coverage (matching effort), not perfect correctness. The correctness of those attempted rules is then picked up progressively downstream in Q2–Q6.
@@
closeComplaint(X), -consent(X), -clausesCDEComplied(X) $\Rightarrow$ [O] closeComplaint(X)
Note that the consequence of this rule also appears as its antecedent. Furthermore, this question checks for syntactically identical rules that appear more than once within a law snippet.
@@
However, there are also more subtle issues filtered by this question, for example, when a LLM combines several aspects with a logical “and”, even though they should be connected with a logical “or” according to the legal text. This question also verifies whether two rules of a formalization convey the same meaning and are therefore redundant, even though they are not syntactically identical.
@@
-consentConsumer(X), -compliedWithClauseC(X),-compliedWithClauseD(X),
-compliedWithClauseE(X) $\Rightarrow$ [F] closeComplaint(X)
@@
Let $l$ be a law snippet. We denote by $\mathcal{R}(l)$ the set of rules generated by the LLM from $l$ and by $\mathcal{G}(l)$ the set of rules from the gold standard corresponding to $l$. We write $r\sim r^{\prime}$ if a rule $r\in\mathcal{R}(l)$ can be meaningfully matched to a rule $r^{\prime}\in\mathcal{G}(l)$ despite possible minor errors as explained earlier. To assess the quality of the generated formalization $\mathcal{R}(l)$, we define a success score $S(l)\in[0,1]$ that captures both the coverage of $\mathcal{G}(l)$ and the correctness of the generated rules in $\mathcal{R}(l)$.
Let $\text{Q1}(l)\in[0,1]$ denote the proportion of rules in $\mathcal{R}(l)$ w.r.t. $\mathcal{G}(l)$ that were attempted to be formalized by the LLM. For instance, if the gold standard snippet contains two rules and the LLM generates two flawed but still alignable rules, we assign $\text{Q1}(l)=1.0$. Conversely, if the LLM generates only one rule that matches one of the two gold standard rules, we assign $\text{Q1}(l)=0.5$. In this way, Q1 enforces that all gold standard rules should be addressed, even if not perfectly realized. Formally, $\text{Q1}(l)$ can be defined as:
$\text{Q1}(l)\;=\;\frac{\bigl|\{\,r\in\mathcal{R}(l)\;\mid\;\exists\,r^{\prime}\in\mathcal{G}(l)\text{ such that }r\sim r^{\prime}\}\bigr|}{|\mathcal{G}(l)|}.$
For each rule $r\in R(l)$, we then evaluate the remaining five binary criteria $\text{Q2}(r),\ldots,\text{Q6}(r)\in\{0,1\}$, asked in sequential order and designed to assess progressively finer aspects of correctness. If any $\text{Qi}(r)$ evaluates to 0, all subsequent questions for that rule are also set to 0.
The success score $S(l)$ for an individual law snippet $l$ is then defined as:
$S(l)=\text{Q1}(l)\times\frac{1}{|\mathcal{R}(l)|}\sum_{r\in\mathcal{R}(l)}\frac{1}{5}\sum_{i=2}^{6}\text{Qi}(r).$
This formulation ensures that Q1 rewards coverage (whether an LLM attempts to capture all gold standard rules), while Q2–Q6 reward the accuracy of those attempts.
@@ 4 Results
In each of the following experiments, we provided LLMs with a prompt containing step-by-step instructions to guide the model through its task. This approach is called Chain-of-Instructions (CoI) prompting [DBLP:conf/jurix/ZinSB24]. Hence, the model is encouraged to solve each subtask step by step until the final answer is reached [DBLP:journals/corr/abs-2402-11532]. This method contrasts with Chain-of-Thoughts (CoT), which usually depends more on implicit reasoning [DBLP:conf/jurix/ZinSB24] – especially for Zero-Shot-CoT, where just a sentence like “Let’s think step by step” is appended to the prompt [DBLP:journals/corr/abs-2401-14295].
Moreover, we use few-shot learning [DBLP:conf/nips/BrownMRSKDNSSAA20], where we provide the LLM with a few examples (input-output pairs) in the prompt to demonstrate how to solve the task.